OpenStudy (anonymous):

If the sum of three consecutive integers is 60, what is the smallest integer?

OpenStudy (anonymous):

@cvliboi

OpenStudy (michele_laino):

hint: \[n + \left( {n + 1} \right) + \left( {n + 2} \right) = 60\] what is n?

OpenStudy (anonymous):

18 or 19

OpenStudy (michele_laino):

18

OpenStudy (anonymous):

im in a test you cant just tell me

OpenStudy (michele_laino):

n=19 n+1=20 n+2=21

OpenStudy (michele_laino):

sorry n=18 is wrong answer

OpenStudy (anonymous):

ok thanxs

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

.